# On the Compatibility of Constructive Predicative Mathematics with Weyl’s Classical Predicativity

### v3

@@ 1 Introduction
The book Das Kontinuum [55, 56] is a milestone in the foundations of mathematics and, in particular, of classical predicative mathematics. Weyl’s monograph is mainly devoted to the predicative reformulation of real analysis.
Another major reformulation of real analysis within a foundation weaker than the usual classical one was given by Bishop in [6]. While Bishop followed Brouwer in adopting intuitionistic logic, he refrained in his approach from including Brouwer’s continuity principles that are inconsistent with classical mathematics. Bishop aimed to develop constructive mathematics not in opposition to classical mathematics but as a generalization of the latter.
After the publication of Bishop’s book, there was a flourishing of foundations for constructive mathematics aimed at formally capturing Bishop’s approach. A relevant example in axiomatic set theory is Aczel’s constructive set theory CZF [1], and in type theory Martin-Löf Type Theory MLTT [39, 40] and the more recent Homotopy Type Theory HoTT  in [53] inspired by Voevodsky’s Univalent Foundations. In particular, HoTT  gained a lot of attention for its application to the formalization of synthetic homotopy theory and higher category theory. However, these systems are sometimes not even mutually compatible and this consideration led the second author in joint work with G. Sambin to conceive the Minimalist Foundation (for short MF) in [34], later fully developed as a two-level system in [24]. The two-level structure of MF in [24], which includes an intensional level, an extensional level, and an interpretation of the latter into the first, was introduced to better achieve the compatibility of MF with constructive and classical foundations including those mentioned (see [24, 9]) as well as internal theories of categorical universes such as that of elementary topoi in [22, 23].
@@
Here we first discuss the source of this incompatibility. The reason lies in their possibility of defining sets by quantifying over (the exponentiation of) functional relations. Such a possibility is not allowed in modern formulations of Weyl’s system [5, 4, 3, 16]. More formally, we show that the incompatibility of CZF, MLTT and HoTT with classical predicativity à la Weyl reduces to the fact of being able to interpret Heyting arithmetic in all finite types with the addition of the internal rule of number-theoretic unique choice, identifying functional relations over natural numbers with a primitive notion of function defined as $\lambda$-terms of type theory.
@@
First of all, the intuitionistic version of the subsystem of second-order arithmetic known as ACA [48] can be interpreted in the extensional level of MF by preserving the meaning of all the logical operators, as well as the classical ACA can be analogously interpreted in the classical extension of MF. Therefore, Weyl’s reconstruction of analysis can be performed in the classical version of MF, since it is known from results of reverse mathematics in [16, 17, 48] that the formalization of Weyl’s analysis can be performed at least in ACA, whilst the proof-theoretic strength of Weyl’s original logical system exceeds it, see [5, 4, 3, 16].
Second, the classical version of MF is equiconsistent, as shown in [27, 33], with the original constructive MF, whose proof-theoretic strength is strictly predicative à la Feferman being interpretable in first-order Martin-Löf’s type theory with one universe [24] and directly in Feferman’s $\widehat{ID_{1}}$ (see [20]).
@@
As a consequence, these limitations call for a point-free constructive development of topology as advocated by Martin-Löf and Sambin with the introduction of Formal Topology [45] to avoid the use of Brouwer’s Fan theorem as explained in [35]. Evidence that such an approach can be successfully developed is provided by the point-free results in [7, 42, 43, 8, 21, 47].
@@
In the future, besides determining the exact proof-theoretic strength of MF, which is currently an open problem, we intend to investigate whether the extensions of MF  with inductive and coinductive definitions in [28, 29, 32] are still equiconsistent with their classical counterparts as established for MF in [27, 33], or at least these classical counterparts are still predicative. A positive answer to one of these questions would provide more expressive foundations than MF capable of formalizing a whole development of point-free topology, including for example the results in [37, 15, 10, 47], while still keeping the compatibility with both Bishop’s constructive mathematics and Weyl’s classical predicative mathematics.
@@ 2 Weyl’s Classical Predicativity
Das Kontinuum [55, 56] can be regarded as the first systematic attempt to develop predicative mathematics. The book’s main focus is the foundation of analysis and the reconstruction of many fundamental results in this area without resorting to impredicative principles. Examples of theorems shown by Weyl in his framework include sequential versions of the Heine-Borel Theorem and the Least Upper Bound principle, Uniform continuity for continuous functions over a closed interval, the existence of a maximum and a minimum for them as well as the Fundamental Theorem of the Calculus.
Therefore, the main merit of Das Kontinuum was to show that predicative mathematics was a viable project. Feferman has shown that the above mentioned results can be formalized in the subsystem of second-order arithmetic known as ACA0 and further evidence in support of this claim has been provided by the Reverse Mathematics program [48]. In this section, we review some basic tenets of Weyl’s approach by relating them to the Minimalist Foundation, which will be introduced later in Section 6.
@@
Weyl’s conception of set bears some resemblance to the way sets are understood in Martin-Löf’s type theory (MLTT) and also in both levels of the Minimalist Foundation (MF), which will be presented in the next sections. In such theories, a set is determined by rules for forming its elements, equipped with an elimination rule that allows proving properties by induction over it and defining functions by recursion from it to another set, see [40].
Furthermore, Weyl considered the notion of set (and function) underlying the classical foundation of analysis to be irremediably circular. Weyl noted that many basic definitions of analysis (e.g. that of least upper bound) were actually circular, and decided to reformulate analysis by avoiding such viciously circular definitions in favour of a predicative development. In particular, such a circularity clearly shows up in Dedekind’s construction of the real numbers: real numbers defined in this way do not form an extensionally determinate collection [^1] See [5, 14] for a detailed analysis of this concept.. Weyl elaborates on this point in a letter to Hölder, published as appendix in [56]:
‘If $A$ is a property of properties, then one may form the property $P_{A}$ which belongs to an object $x$ if and only if there is a property constructed by means of these principles which belongs to $x$ and itself possesses the property $A$. That would be a blatant circulus vitiosus; yet our current version of analysis commits this error and I consider it ground for censure.’ ([56], p.113)
@@
In the Minimalist Foundation, this distinction between sets and collections is explicitly assumed. Accordingly, in Weyl’s work collections exceed the power-collection of a set, as shown in [4, 3, 5], but they are not allowed to enter into the definition of new sets.
A major distinction between sets and collections in the Minimalist Foundation is that, whilst both entities are generated by rules forming their canonical elements as sets in Martin-Löf’s type theory, sets distinguish from collections for being equipped with an induction principle and a recursive definition of functions from them toward any other entity. Collections instead do not enjoy any induction principle or recursive definition from them, since inductive or recursive definitions acting on them would not be invariant under extensions of the theory with new sets. This approach aligns closely with the idea that predicative sets should be invariant, which traces back to Poincaré and has been carefully analyzed in [13].
@@
Finally, we want to point out that Weyl in Das Kontinuum places a significant emphasis on his notion of function and contrasts it with the traditional set-theoretical definition (see [56], pp.33-34). According to Avron’s analysis in [5], for Weyl, a function is always determined by a rule. In other words, a function cannot be given by an arbitrary relation satisfying the functionality condition, as in set theory; rather, there are syntactic constraints on the definability of the relation that must be respected.
@@
Weyl’s insights about functions as rule-based objects align, in some way, with the fact that the Minimalist Foundation distinguishes the notion of function as functional relation from that of function as a primitive operation defined by a $\lambda$-term, resembling the Brouwer’s notion of lawlike sequence (see [46]). The distinct character of these two notions of function is a peculiarity of MF among foundations for constructive predicative mathematics. While it might be argued that Weyl foreshadowed this distinction, nevertheless his technical treatment of the notion of function is not exactly the same.
There have been various attempts in the literature to formalize Weyl’s system in a modern fashion, since Weyl’s original presentation presents some ambiguities. An influential reconstruction of Weyl’s system has been carried out by Feferman in several works [16, 17, 18]. Feferman has extensively argued that Weyl’s system should correspond to the theory $\mathbf{ACA}_{0}$ or, if the full induction scheme is accepted, to $\mathbf{ACA}$, which are well-known subsystems of second-order arithmetic [48]. But such a reconstruction has been criticized by Adams and Luo in [4, 3], and more recently by Avron in [5]. They observed that in certain sections of Das Kontinuum, Weyl introduces non-arithmetical sets and makes use of induction for non-arithmetical formulas. Therefore, they concluded that Weyl’s system should be stronger than Feferman’s reformulation of Weyl’s system, even if ${\bf ACA}_{0}$ should be sufficient to reconstruct predicative analysis as developed in Das Kontinuum.
The systems introduced in the above-mentioned papers are many-sorted logics axiomatized over a dependent type theory, in a way similar to both levels of the Minimalist Foundation. They all differ in the way exponentiation of functions is treated. Indeed, Adams and Luo’s system in [4, 3] includes some exponentiation of $\lambda$-functions in the form of function types, which are treated as large types in the sense that they are not contained in the universe of small types, that in turn can be thought of as the universe of sets. Instead, Avron in [5] restricts the possibility of forming exponentials by assuming that only certain types might serve as codomains by not allowing to define sets by quantification over functions. On the contrary, in MF there is no exponentiation of functions between collections but only sets exponentiating $\lambda$-functions between sets.
@@ 3 Constructive Predicative Theories
Among the most popular frameworks introduced as foundations for Bishop’s constructive mathematics there are Constructive Zermelo-Fraenkel set theory $\mathbf{CZF}$ [1] and Martin-Löf Type Theory MLTT  [40]. More recently, an extension of the latter called Homotopy Type Theory HoTT [53] has gained much attention, especially for its applications to synthetic homotopy theory and higher categories. All these systems are examples of constructive and predicative theories. However, the predicativity of these theories is closely tied to the adoption of a constructive logic. Indeed, whenever they are extended with the law of excluded middle or equivalent classical principles, they become impredicative and thus they are not compatible with classical predicativity à la Weyl. Briefly, the issue is that such theories are closed under the exponentiation of functional relations and the two-element set that, when classical logic is added, are strong enough to allow for impredicative constructions. Therefore, we cannot establish the predicativity of their classical counterparts, as it happens instead in the case of Heyting arithmetic or the Minimalist Foundation, which we will introduce in detail in Section 6.
@@ 3.1 Aczel’s Constructive Zermelo-Fraenkel set theory
CZF [1] is a theory formulated in the same language of classical ZF and based on intuitionistic logic. The theory includes the following axioms: extensionality, pairing, union, restricted separation, infinity, $\in$-induction, strong collection, and subset collection.
@@
In [2], it is shown that this proof can be carried out in a weaker subsystem of CZF. On the other hand, the weaker principles of countable choice and dependent choice can be added to CZF without leading to undesirable results.
@@ 3.2 Martin-Löf’s type theory
Martin-Löf Type Theory (MLTT) [40] is a foundation for constructive mathematics, which was explicitly introduced to formalize Bishop’s mathematics [38]. The theory is not presented in the usual language of first-order logic, but it is formulated in the language of dependent type theory.
@@
Henceforth, we can refer in MLTT  with the name function to either a functional relation or a $\lambda$-function without ambiguity. As anticipated in section 2, the two notions of function are instead kept distinct in the Minimalist Foundation as well as in some modern presentations of Weyl’s system as those given in [3, 5].
Due to the absence of quotient sets in MLTT, real numbers, both à la Dedekind or à la Cauchy, are formally represented there using setoids whose elements are then indexed by a set. A setoid [41], called set by Bishop in ([6], p. 13), is a pair $(A,=_{A})$, consisting of a set called support equipped with a suitable equivalence relation $a=_{A}a^{\prime}\ [a,a^{\prime}:A]$
For example, Bishop reals in [6] can be represented as the setoid whose support is given by the set of Cauchy regular sequences of rational numbers equipped with the equivalence relation $=_{\mathcal{R_{C}}}$ as in ([6], pp.15-16). Bishop defined them by using the notion of operation which we can identify in MLTT  with the notion of $\lambda$-function, i.e. with a $\lambda$-term of MLTT. Moreover, for what we have said about the notion of function in MLTT, Cauchy reals defined as Cauchy functional relations can be identified in MLTT with Bishop reals. For unifying the various concepts, we name them Cauchy sequences à la Bishop.
@@
A Dedekind cut on the rationals is a pair $(L,U)$ of subsets $L,U\subseteq\mathbb{Q}$ that satisfy the usual conditions of being inhabited, disjoint, open, monotone, and located, all defined as in [36].
Notice that the subsets $L,U$ are defined as elements of the setoid of propositional functions $L,U:\mathbb{Q}\rightarrow\mathcal{U}_{k}$ for some universe $\mathcal{U}_{k}$ quotiented under equiprovability. Therefore, ‘$a\ \varepsilon\ L$’ means that there exists a proof-term of $L(a)$.
@@
Given a Dedekind cut as in def. 3.5 in $\mathbb{R}_{D_{k}}$ for any natural number $k$, by using the axiom of choice we can define a Bishop real given by an operation $\lambda n.x_{n}$, usually written $(x_{n})_{n}$, as in Prop. 17 in [7] or section 4.3 in [35], whose cut $(L_{0},D_{0})$ in $\mathbb{R}_{D_{0}}$ is defined as
@@
and is propositionally equivalent to the starting cut $(L,U)$ in $\mathbb{R}_{D_{k}}$ thanks to the correspondence between Dedekind cuts, formal points of point-free topology of real numbers and Bishop reals in section 4.3 of [35]. ∎

### v4

@@ 1 Introduction
The book Das Kontinuum [Weyl+1918, weyl1994continuum] is a milestone in the foundations of mathematics and, in particular, of classical predicative mathematics. Weyl’s monograph is mainly devoted to the predicative reformulation of real analysis.
Another major reformulation of real analysis within a foundation weaker than the usual classical one was given by Bishop in [bishop1967]. While Bishop followed Brouwer in adopting intuitionistic logic, he refrained in his approach from including Brouwer’s continuity principles that are inconsistent with classical mathematics. Bishop aimed to develop constructive mathematics not in opposition to classical mathematics but as a generalization of the latter.
After the publication of Bishop’s book, there was a flourishing of foundations for constructive mathematics aimed at formally capturing Bishop’s approach. A relevant example in axiomatic set theory is Aczel’s constructive set theory CZF [aczel78], and in type theory Martin-Löf Type Theory MLTT [martin-lof:bibliopolis, nordstrom1990programming] and the more recent Homotopy Type Theory HoTT in [hottbook] inspired by Voevodsky’s Univalent Foundations. In particular, HoTT gained a lot of attention for its application to the formalization of synthetic homotopy theory and higher category theory. However, these systems are sometimes not even mutually compatible and this consideration led the second author in joint work with G. Sambin to conceive the Minimalist Foundation (for short MF) in [ms05], later fully developed as a two-level system in [m09]. The two-level structure of MF in [m09], which includes an intensional level, an extensional level, and an interpretation of the latter into the first, was introduced to better achieve the compatibility of MF with constructive and classical foundations including those mentioned (see [m09, cm2024]) as well as internal theories of categorical universes such as that of elementary topoi in [LS1986, Maietti05].
@@
Here we first discuss the source of this incompatibility. The reason lies in their possibility of defining sets by quantifying over (the exponentiation of) functional relations. Such a possibility is not allowed in modern formulations of Weyl’s system [DBLP:journals/bsl/Avron20, adams2010weyl, DBLP:journals/apal/AdamsL10, feferman1988weyl]. More formally, we show that the incompatibility of CZF, MLTT and HoTT with classical predicativity à la Weyl reduces to the fact of being able to interpret Heyting arithmetic in all finite types with the addition of the internal rule of number-theoretic unique choice, identifying functional relations over natural numbers with a primitive notion of function defined as $\lambda$-terms of type theory.
@@
First of all, the intuitionistic version of the subsystem of second-order arithmetic known as ACA [simpson2009subsystems] can be interpreted in the extensional level of MF by preserving the meaning of all the logical operators, as well as the classical ACA can be analogously interpreted in the classical extension of MF. Therefore, Weyl’s reconstruction of analysis can be performed in the classical version of MF, since it is known from results of reverse mathematics in [feferman1988weyl, feferman1998light, simpson2009subsystems] that the formalization of Weyl’s analysis can be performed at least in ACA, whilst the proof-theoretic strength of Weyl’s original logical system exceeds it, see [DBLP:journals/bsl/Avron20, adams2010weyl, DBLP:journals/apal/AdamsL10, feferman1988weyl].
Second, the classical version of MF is equiconsistent, as shown in [talkmillyober, MSdoubleneg], with the original constructive MF, whose proof-theoretic strength is strictly predicative à la Feferman being interpretable in first-order Martin-Löf’s type theory with one universe [m09] and directly in Feferman’s $\widehat{ID_{1}}$ (see [IMMS]).
@@
As a consequence, these limitations call for a point-free constructive development of topology as advocated by Martin-Löf and Sambin with the introduction of Formal Topology [sambin1987] to avoid the use of Brouwer’s Fan theorem as explained in [MSpointfree]. Evidence that such an approach can be successfully developed is provided by the point-free results in [CN95, Palmgren05, pal07, spa, kawai2023, sambinbook].
@@
In the future, besides determining the exact proof-theoretic strength of MF, which is currently an open problem, we intend to investigate whether the extensions of MF with inductive and coinductive definitions in [MMR21, MMR22, MS2023] are still equiconsistent with their classical counterparts as established for MF in [talkmillyober, MSdoubleneg], or at least these classical counterparts are still predicative. A positive answer to one of these questions would provide more expressive foundations than MF capable of formalizing a whole development of point-free topology, including for example the results in [mv04, curicind, CPS11, sambinbook], while still keeping the compatibility with both Bishop’s constructive mathematics and Weyl’s classical predicative mathematics.
@@ 2 Weyl’s Classical Predicativity
Das Kontinuum [Weyl+1918, weyl1994continuum] can be regarded as the first systematic attempt to develop predicative mathematics. The book’s main focus is the foundation of analysis and the reconstruction of many fundamental results in this area without resorting to impredicative principles. Examples of theorems shown by Weyl in his framework include sequential versions of the Heine-Borel Theorem and the Least Upper Bound principle, Uniform continuity for continuous functions over a closed interval, the existence of a maximum and a minimum for them as well as the Fundamental Theorem of the Calculus.
Therefore, the main merit of Das Kontinuum was to show that predicative mathematics was a viable project. Feferman has shown that the above mentioned results can be formalized in the subsystem of second-order arithmetic known as ACA0 and further evidence in support of this claim has been provided by the Reverse Mathematics program [simpson2009subsystems]. In this section, we review some basic tenets of Weyl’s approach by relating them to the Minimalist Foundation, which will be introduced later in Section 6.
@@
Weyl’s conception of set bears some resemblance to the way sets are understood in Martin-Löf’s type theory (MLTT) and also in both levels of the Minimalist Foundation (MF), which will be presented in the next sections. In such theories, a set is determined by rules for forming its elements, equipped with an elimination rule that allows proving properties by induction over it and defining functions by recursion from it to another set, see [nordstrom1990programming].
Furthermore, Weyl considered the notion of set (and function) underlying the classical foundation of analysis to be irremediably circular. Weyl noted that many basic definitions of analysis (e.g. that of least upper bound) were actually circular, and decided to reformulate analysis by avoiding such viciously circular definitions in favour of a predicative development. In particular, such a circularity clearly shows up in Dedekind’s construction of the real numbers: real numbers defined in this way do not form an extensionally determinate collection [^1] See [DBLP:journals/bsl/Avron20, crosilla2023weyl] for a detailed analysis of this concept.. Weyl elaborates on this point in a letter to Hölder, published as appendix in [weyl1994continuum]:
‘If $A$ is a property of properties, then one may form the property $P_{A}$ which belongs to an object $x$ if and only if there is a property constructed by means of these principles which belongs to $x$ and itself possesses the property $A$. That would be a blatant circulus vitiosus; yet our current version of analysis commits this error and I consider it ground for censure.’ ([weyl1994continuum], p.113)
@@
In the Minimalist Foundation, this distinction between sets and collections is explicitly assumed. Accordingly, in Weyl’s work collections exceed the power-collection of a set, as shown in [adams2010weyl, DBLP:journals/apal/AdamsL10, DBLP:journals/bsl/Avron20], but they are not allowed to enter into the definition of new sets.
A major distinction between sets and collections in the Minimalist Foundation is that, whilst both entities are generated by rules forming their canonical elements as sets in Martin-Löf’s type theory, sets distinguish from collections for being equipped with an induction principle and a recursive definition of functions from them toward any other entity. Collections instead do not enjoy any induction principle or recursive definition from them, since inductive or recursive definitions acting on them would not be invariant under extensions of the theory with new sets. This approach aligns closely with the idea that predicative sets should be invariant, which traces back to Poincaré and has been carefully analyzed in [Crosilla2022].
@@
Finally, we want to point out that Weyl in Das Kontinuum places a significant emphasis on his notion of function and contrasts it with the traditional set-theoretical definition (see [weyl1994continuum], pp.33-34). According to Avron’s analysis in [DBLP:journals/bsl/Avron20], for Weyl, a function is always determined by a rule. In other words, a function cannot be given by an arbitrary relation satisfying the functionality condition, as in set theory; rather, there are syntactic constraints on the definability of the relation that must be respected.
@@
Weyl’s insights about functions as rule-based objects align, in some way, with the fact that the Minimalist Foundation distinguishes the notion of function as functional relation from that of function as a primitive operation defined by a $\lambda$-term, resembling the Brouwer’s notion of lawlike sequence (see [sambin2008continuity]). The distinct character of these two notions of function is a peculiarity of MF among foundations for constructive predicative mathematics. While it might be argued that Weyl foreshadowed this distinction, nevertheless his technical treatment of the notion of function is not exactly the same.
There have been various attempts in the literature to formalize Weyl’s system in a modern fashion, since Weyl’s original presentation presents some ambiguities. An influential reconstruction of Weyl’s system has been carried out by Feferman in several works [feferman1988weyl, feferman1998light, feferman2000significance]. Feferman has extensively argued that Weyl’s system should correspond to the theory $\mathbf{ACA}_{0}$ or, if the full induction scheme is accepted, to $\mathbf{ACA}$, which are well-known subsystems of second-order arithmetic [simpson2009subsystems]. But such a reconstruction has been criticized by Adams and Luo in [adams2010weyl, DBLP:journals/apal/AdamsL10], and more recently by Avron in [DBLP:journals/bsl/Avron20]. They observed that in certain sections of Das Kontinuum, Weyl introduces non-arithmetical sets and makes use of induction for non-arithmetical formulas. Therefore, they concluded that Weyl’s system should be stronger than Feferman’s reformulation of Weyl’s system, even if ${\bf ACA}_{0}$ should be sufficient to reconstruct predicative analysis as developed in Das Kontinuum.
The systems introduced in the above-mentioned papers are many-sorted logics axiomatized over a dependent type theory, in a way similar to both levels of the Minimalist Foundation. They all differ in the way exponentiation of functions is treated. Indeed, Adams and Luo’s system in [adams2010weyl, DBLP:journals/apal/AdamsL10] includes some exponentiation of $\lambda$-functions in the form of function types, which are treated as large types in the sense that they are not contained in the universe of small types, that in turn can be thought of as the universe of sets. Instead, Avron in [DBLP:journals/bsl/Avron20] restricts the possibility of forming exponentials by assuming that only certain types might serve as codomains by not allowing to define sets by quantification over functions. On the contrary, in MF there is no exponentiation of functions between collections but only sets exponentiating $\lambda$-functions between sets.
@@ 3 Constructive Predicative Theories
Among the most popular frameworks introduced as foundations for Bishop’s constructive mathematics there are Constructive Zermelo-Fraenkel set theory $\mathbf{CZF}$ [aczel78] and Martin-Löf Type Theory MLTT [nordstrom1990programming]. More recently, an extension of the latter called Homotopy Type Theory HoTT [hottbook] has gained much attention, especially for its applications to synthetic homotopy theory and higher categories. All these systems are examples of constructive and predicative theories. However, the predicativity of these theories is closely tied to the adoption of a constructive logic. Indeed, whenever they are extended with the law of excluded middle or equivalent classical principles, they become impredicative and thus they are not compatible with classical predicativity à la Weyl. Briefly, the issue is that such theories are closed under the exponentiation of functional relations and the two-element set that, when classical logic is added, are strong enough to allow for impredicative constructions. Therefore, we cannot establish the predicativity of their classical counterparts, as it happens instead in the case of Heyting arithmetic or the Minimalist Foundation, which we will introduce in detail in Section 6.
@@ 3.1 Aczel’s Constructive Zermelo-Fraenkel set theory
CZF [aczel78] is a theory formulated in the same language of classical ZF and based on intuitionistic logic. The theory includes the following axioms: extensionality, pairing, union, restricted separation, infinity, $\in$-induction, strong collection, and subset collection.
@@
In [AczelRath], it is shown that this proof can be carried out in a weaker subsystem of CZF. On the other hand, the weaker principles of countable choice and dependent choice can be added to CZF without leading to undesirable results.
@@ 3.2 Martin-Löf’s type theory
Martin-Löf Type Theory (MLTT) [nordstrom1990programming] is a foundation for constructive mathematics, which was explicitly introduced to formalize Bishop’s mathematics [ml1975]. The theory is not presented in the usual language of first-order logic, but it is formulated in the language of dependent type theory.
@@
Henceforth, we can refer in MLTT with the name function to either a functional relation or a $\lambda$-function without ambiguity. As anticipated in section 2, the two notions of function are instead kept distinct in the Minimalist Foundation as well as in some modern presentations of Weyl’s system as those given in [DBLP:journals/apal/AdamsL10, DBLP:journals/bsl/Avron20].
Due to the absence of quotient sets in MLTT, real numbers, both à la Dedekind or à la Cauchy, are formally represented there using setoids whose elements are then indexed by a set. A setoid [palmgren2005bishop], called set by Bishop in ([bishop1967], p. 13), is a pair $(A,=_{A})$, consisting of a set called support equipped with a suitable equivalence relation $a=_{A}a^{\prime}\ [a,a^{\prime}:A]$ .
For example, Bishop reals in [bishop1967] can be represented as the setoid whose support is given by the set of Cauchy regular sequences of rational numbers equipped with the equivalence relation $_{\mathcal{R_{C}}}$ as in ([bishop1967], pp.15-16). Bishop defined them by using the notion of operation which we can identify in MLTT with the notion of $\lambda$-function, i.e. with a $\lambda$-term of MLTT. Moreover, for what we have said about the notion of function in MLTT, Cauchy reals defined as Cauchy functional relations can be identified in MLTT with Bishop reals. For unifying the various concepts, we name them Cauchy sequences à la Bishop.
@@
A Dedekind cut on the rationals is a pair $(L,U)$ of subsets $L,U\subseteq\mathbb{Q}$ that satisfy the usual conditions of being inhabited, disjoint, open, monotone, and located, all defined as in [maiettisambinhand].
Notice that the subsets $L,U$ are defined as elements of the setoid of propositional functions $L,U:\mathbb{Q}\rightarrow\mathcal{U}_{k}$ for some universe $\mathcal{U}_{k}$ quotiented under equiprovability. Therefore, $a\ \varepsilon\ L$’ means that there exists a proof-term of $L(a)$.
@@
Given a Dedekind cut as in def. 3.5 in $\mathbb{R}_{D_{k}}$ for any natural number $k$, by using the axiom of choice we can define a Bishop real given by an operation $\lambda n.x_{n}$, usually written $(x_{n})_{n}$, as in Prop. 17 in [CN95] or section 4.3 in [MSpointfree], whose cut $(L_{0},D_{0})$ in $\mathbb{R}_{D_{0}}$ is defined as
@@
and is propositionally equivalent to the starting cut $(L,U)$ in $\mathbb{R}_{D_{k}}$ thanks to the correspondence between Dedekind cuts, formal points of point-free topology of real numbers and Bishop reals in section 4.3 of [MSpointfree]. ∎